The perimeter of an isosceles triangle is 22m. Two of the sides are the same length, and the third side is 2m less than the length of the other sides. What are the lengths of all sides of the triangle?
Answer by jim_thompson5910(35256)

Let x = length of equal sides, so x-2 is the third side.

Add up all the sides and set them equal to the perimeter 22. Then solve for x.

x + x + (x-2) = 22

3x - 2 = 22

3x = 22+2

3x = 24

x = 24/3

x = 8

x-2 = 8-2 = 6

So the two equal sides are 8 meters and the third side is 6 meters.

Check

P = S1 + S2 + S3

22 = 8 + 8 + 6

22 = 16 + 6

22 = 22

Answer checks out